OpenStudy (anonymous):

how do you factor 2x^2=6x+20

OpenStudy (zzr0ck3r):

2x^2-6x-20 = 0

OpenStudy (zzr0ck3r):

2(x^2-3x-10) = 0 2(x^2+2x-5x-10) = 0 can you do the rest?